# CDQuant: Accurate Post-training Weight Quantization of Large Pre-trained Models using Greedy Coordinate Descent

### v2

@@ 4 Experiments
Quantization.
In our experiments, we focus on weight-only quantization. We present results for two scenarios: (1) quantizing the FFN layers and (2) quantizing both the FFN and attention weights. In the former setting, all attention layers are quantized to INT8 using MinMax quantization, and low-bit quantization schemes are only applied to the feed-forward layers. It’s worth pointing out that this setting is very relevant in practice because FFN layers account for majority of the latency, and quantizing attention weights yields negligible latency reduction at the expense of model quality (Samaga et al., 2024). We perform both per-channel and sub-channel quantization (with a group size of $128$). We mainly focus on INT2, INT3 and INT4 quantization.
Models.
We use PaLM2 models (Anil et al., 2023) in all our experiments. In particular, we use PaLM2-Gecko, PaLM2-Otter, PaLM2-Bison models. These models vary in size, with PaLM2-Gecko being the smallest, followed by PaLM2-Otter, and finally PaLM2-Bison.
Baselines.
@@
Since our primary goal is to demonstrate that CDQuant gets improved performance over GPTQ, we have chosen it as the primary baseline in most of our experiments. To further demonstrate the our technique can be used as a plug-and-play replacement for GPTQ, we also include comparisons with AWQ (Lin et al., 2023). For all our experiments, we initialize GPTQ with OWC, and run GPTQ for $T=d_{\text{in}}$ steps. To ensure stability and generalization, GPTQ regularizes its Hessian matrix by adding a scaled identity matrix ($\lambda I$). Tuning this $\lambda$ for every (model, layer) pair is infeasible. So, we determine a single optimal value using the PaLM2-Gecko model, and apply it universally.
CDQuant.
@@
We evaluate both the coordinate descent variants described in Section 3.1: CD (Algorithm 1), BCD (Algorithm 2). For per-channel quantization, we initialize CD with Optimal Weight Clipping (OWC), and for sub-channel quantization, we additionally include initialization with OWC-CD. BCD is always initialized with CD in our experiments. Unless otherwise stated, both CD and BCD are run for $T=d_{\text{in}}$ iterations, OWC-CD is run for $d_{\text{in}}/g$ iterations, where $g$ is the group size. Similar to GPTQ, we regularize the Hessian matrix used in CD, BCD by adding $\lambda I$. We determine a reasonable value for $\lambda$ using the PaLM2-Gecko model, and use it in all our experiments.
Evaluation.
@@
Following recent works (Frantar et al., 2022; Ma et al., 2024), we evaluate all algorithms using two key metrics: perplexity and downstream task performance. To ensure a robust evaluation, we calculate perplexity using a $100$ million token subset derived from the PaLM2 training mixture. We chose to do this since the evaluation sets for PTB (Marcus et al., 1994) and WikiText2 (Merity et al., 2017) datasets, that are commonly used in the field, are orders of magnitude smaller (83K and 246K tokens in their respective test sets, and 512K tokens subsampled from the C4 validation set  (Frantar et al., 2022)), and thus can have high variance, and may not accurately reflect the model’s true performance. We use TriviaQA (Joshi et al., 2017), SQuAD (Rajpurkar et al., 2018), NaturalQuestions (Kwiatkowski et al., 2019) and WebQuestions (Berant et al., 2013) to evaluate generation capabilities of the quantized models, and to evaluate their reasoning capabilities, we test on ARC-c, ARC-e (Clark et al., 2018), HellaSwag (Zellers et al., 2019), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020) and WinoGrande (Sakaguchi et al., 2020). For downstream evaluations, we consider both zero-shot, and one-shot settings. Finally, to determine which optimization technique is most effective at solving the layer-wise objective in Equation (1), we evaluate the final solution’s objective value using the same $100$ million tokens that we used to compute perplexity.
Training.
@@
All techniques are calibrated using $1280$ data points, where each data point has $2048$ tokens. For OWC, we use a grid size of $50$ to find the most optimal $\gamma$. We used 8 Nvidia H100100100100 GPUs for quantizing the models.
@@
Table 1: Table presents pretraining perplexity numbers for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights. $wx,ay,gz$ in the config column corresponds to $x$-bit weights, $y$-bit activations and group/sub-channel size of $z$.
Config Method PaLM2-Gecko PaLM2-Otter PaLM2-Bison
w16a16 - 7.948 5.984 5.298
w3a16 OWC $12.570$ $17.928$ $6.169$
GPTQ $11.347$ $7.176$ $5.774$
CD $10.920$ $7.002$ $5.739$
BCD(k=2) $\mathbf{10.898}$ $\mathbf{6.979}$ $\mathbf{5.733}$
w3a16g128 OWC $11.597$ $8.342$ $5.847$
GPTQ $10.414$ $6.635$ $5.677$
CD $10.273$ $6.655$ $5.656$
BCD(k=2) $10.259$ $6.545$ $5.654$
OWC-CD $10.706$ $6.635$ $5.686$
OWC-CD + CD $10.143$ $6.528$ $5.650$
OWC-CD + BCD(k=2) $\mathbf{10.138}$ $\mathbf{6.527}$ $\mathbf{5.647}$
w4a16 OWC $8.946$ $6.693$ $5.475$
GPTQ $8.764$ $6.249$ $5.417$
CD $8.694$ $6.195$ $5.407$
BCD(k=2) $\mathbf{8.691}$ $\mathbf{6.192}$ $\mathbf{5.405}$
w4a16g128 OWC $8.613$ $6.264$ $5.401$
GPTQ $8.498$ $6.112$ $5.377$
CD $8.456$ $6.097$ $5.373$
BCD(k=2) $8.454$ $6.097$ $5.372$
OWC-CD $8.519$ $6.106$ $5.377$
OWC-CD + CD $8.436$ $6.092$ $5.371$
OWC-CD + BCD(k=2) $\mathbf{8.434}$ $\mathbf{6.091}$ $\mathbf{5.371}$
@@
Table 2: Table presents downstream evaluation (one-shot) numbers for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights. Gen, Rank columns correspond to generation and ranking tasks.
@@
w16a16 - $25.85$ $61.44$ $47.2$ $44.31$ $75.43$ $62.98$ $51.08$ $78.28$ $68.00$
w3a16 OWC $18.95$ $55.74$ $41.02$ $21.02$ $65.78$ $47.87$ $47.52$ $77.76$ $65.66$
GPTQ $19.42$ $57.71$ $42.39$ $37.26$ $73.95$ $59.27$ $48.55$ $78.58$ $66.57$
CD $20.69$ $58.05$ $43.11$ $38.12$ $73.37$ $59.27$ $48.47$ $78.56$ $66.53$
BCD(k=2) $20.52$ $58.44$ $43.27$ $38.53$ $73.46$ $59.49$ $48.08$ $78.66$ $66.43$
w3a16g128 OWC $22.83$ $57.02$ $43.35$ $33.18$ $70.81$ $55.76$ $48.08$ $78.34$ $66.24$
GPTQ $22.24$ $58.66$ $44.09$ $40.30$ $74.12$ $60.59$ $49.15$ $78.58$ $66.81$
CD $22.79$ $58.81$ $44.40$ $41.33$ $74.27$ $61.10$ $49.17$ $78.49$ $66.76$
BCD(k=2) $22.64$ $58.85$ $44.37$ $41.19$ $73.94$ $60.84$ $49.02$ $78.51$ $66.71$
OWC-CD $22.21$ $58.69$ $44.10$ $40.84$ $74.33$ $60.93$ $48.54$ $78.29$ $66.39$
OWC-CD + CD $22.44$ $58.84$ $44.28$ $40.87$ $74.41$ $61.00$ $49.22$ $78.86$ $67.00$
OWC-CD + BCD(k=2) $22.15$ $58.77$ $44.13$ $41.19$ $74.65$ $61.27$ $49.08$ $78.3$ $66.61$
w4a16 OWC $23.27$ $60.22$ $45.44$ $41.70$ $74.21$ $61.21$ $50.05$ $78.72$ $67.25$
GPTQ $24.51$ $60.72$ $46.24$ $43.00$ $74.86$ $62.12$ $50.28$ $79.15$ $67.6$
CD $24.60$ $60.62$ $46.22$ $43.21$ $75.17$ $62.39$ $50.09$ $79.21$ $67.56$
BCD(k=2) $24.52$ $60.62$ $46.18$ $43.16$ $75.50$ $62.56$ $50.25$ $79.14$ $67.59$
w4a16g128 OWC $24.63$ $61.27$ $46.61$ $42.31$ $74.76$ $61.78$ $50.49$ $79.38$ $67.82$
GPTQ $25.40$ $61.13$ $46.84$ $43.54$ $75.02$ $62.43$ $50.63$ $79.21$ $67.78$
CD $25.18$ $61.42$ $46.92$ $43.78$ $74.93$ $62.47$ $50.73$ $79.30$ $67.87$
BCD(k=2) $24.95$ $61.49$ $46.87$ $43.77$ $75.04$ $62.53$ $50.84$ $79.56$ $68.07$
OWC-CD $24.59$ $61.40$ $46.67$ $43.67$ $75.31$ $62.65$ $50.65$ $79.04$ $67.69$
OWC-CD + CD $25.12$ $61.44$ $46.91$ $44.00$ $75.27$ $62.76$ $50.62$ $79.21$ $67.78$
OWC-CD + BCD(k=2) $25.00$ $61.50$ $46.90$ $43.75$ $75.20$ $62.62$ $50.64$ $79.17$ $67.76$
Results.
Table 1 presents the perplexity numbers for different quantization techniques applied to FFN layers. It can be seen that both our CD and BCD methods have a clear advantage over GPTQ, leading to lower perplexity scores for all models and quantization levels. The difference is more pronounced for lower bit quantization. This is further highlighted in Table 5, which focuses specifically on INT2 quantization. Here, we see almost $10\%$ improvement in perplexity over GPTQ. We also present (1-shot) downstream evals in Tables 25 (for 0-shot results, and a detailed breakdown of results for each dataset, please refer to Appendix D). In Table 5 we see $2$-$2.5\%$ relative improvement in downstream evals over GPTQ. Furthermore, Table 2 demonstrates that our methods consistently match or exceed GPTQ’s performance across all the settings, with the most substantial improvements observed in low-bit quantization, and smaller models. Finally, we also observe that coordinate descent techniques are better at optimizing the layer-wise objective in Equation (1), than GPTQ. For instance, the average objective value (relative to all 00’s solution) for the GPTQ solution for the $1^{\text{st}}$ feed-forward layer is $0.164$, whereas for CD it is $0.158$, and for BCD(k=$2$) it is $0.157$.
Next, observe that Table 5 also investigates the effect of multiple epochs of CD and BCD, where each epoch corresponds to $d_{\text{in}}$ iterations. For BCD with $k=2$, additional epochs enhance performance. However, for BCD with $k=4$, multiple epochs appear to lead to overfitting, indicating a need for stronger regularization.
FFN+Attention Quantization. In addition to FFN quantization, we also perform FFN + attention weight quantization. However, the inputs to the attention layer are often aligned in a handful of directions. Consequently, performing quantization using such a data leads to a huge drop in performance, as the algorithms would primarily focus on a few directions and ignore the rest. To mitigate this, we clip the largest eigenvalues of the Hessian to ensure a more balanced Hessian. This technique, reminiscent of the weight clipping in OmniQuant (Shao et al., 2023), improves the performance of both GPTQ and CDQuant [^1] One could also rely on existing techniques such as AWQ and SmoothQuant to reduce the effect of outliers. But in our experiments we noticed that both AWQ and SmoothQuant performed poorly compared to the simple eigenvalue clipping technique.. For instance, for PaLM2-Gecko, the perplexity for w3a16-GPTQ improves from $34.872$ to $24.054$ with clipping, and for w4a16-GPTQ, it improves from $10.966$ to $10.005$. Table 4 presents the results from this experiment, where we run all the algorithms on the clipped Hessian. We once again notice that in almost all the settings, our coordinate descent algorithms outperform GPTQ.
Table 3: Perplexity numbers for GPTQ, CD, BCD for quantization of FFN layers of PaLM2-Otter, with and without (w/o) AWQ.
Method PaLM2-Otter
w3a16 w3a16g128
with AWQ w/o AWQ with AWQ w/o AWQ
OWC $22.661$ $17.928$ $6.813$ $8.342$
GPTQ $7.381$ $7.176$ $6.519$ $6.635$
CD $7.244$ $7.002$ $6.475$ $6.655$
BCD(k=2) $\mathbf{7.211}$ $\mathbf{6.979}$ $\mathbf{6.474}$ $\mathbf{6.545}$

### v3

@@ 4 Experiments
In this section, we first establish that CDQuant achieves superior quantization quality compared to GPTQ. Subsequently, we show its versatility by successfully integrating it as a drop-in replacement for GPTQ within existing PTQ methods such as AWQ (Lin et al., 2023), QuIP (Chee et al., 2024), FrameQuant (Adepu et al., 2024).
4.1 Comparison with GPTQ
We first present the experimental setup and then move on to our results. In all the experiments in this section, the attention layers are quantized to INT8 using MinMax quantization, and low-bit quantization schemes are only applied to the feed-forward layers (FFN). See Appendix F for experiments with quantized attention and FFN layers.
Models. Our experiments leverage two families of language models: the open-source Gemma models (Mesnard et al., 2024) and the proprietary PaLM 2 models (Anil et al., 2023). Specifically, we use Gemma-1 7B, Gemma-2 9B, and Gemma-2 27B from the Gemma family, and PaLM2-Gecko, PaLM2-Otter, and PaLM2-Bison from the PaLM2 family. Within the PaLM 2 family, the models increase in size from Gecko to Otter to Bison.
@@ 4.1 Comparison with GPTQ
Baselines. Since our primary goal is to demonstrate that CDQuant gets improved performance over GPTQ, we have chosen it as the primary baseline in most of our experiments in this section. For all our experiments, we initialize GPTQ with OWC, and run GPTQ for $T=d_{\text{in}}$ steps. To ensure stability and generalization, GPTQ regularizes its Hessian matrix by adding a scaled identity matrix ($\lambda I$). Tuning this $\lambda$ for every (model, layer) pair is infeasible. So, we determine a single optimal value using the PaLM2-Gecko model, and apply it universally.
@@
CDQuant. We evaluate both the coordinate descent variants described in Section 3.1: CD (Algorithm 1), BCD (Algorithm 2). For per-channel quantization, we initialize CD with Optimal Weight Clipping (OWC), and for sub-channel quantization, we additionally include initialization with OWC-CD. BCD is always initialized with CD in our experiments. Unless otherwise stated, both CD and BCD are run for $T=d_{\text{in}}$ iterations, OWC-CD is run for $d_{\text{in}}/g$ iterations, where $g$ is the group size. Similar to GPTQ, we regularize the Hessian matrix used in CD, BCD by adding $\lambda I$. We determine a reasonable value for $\lambda$ using the PaLM2-Gecko model, and use it in all our experiments.
@@
Evaluation. Following recent works (Frantar et al., 2022; Ma et al., 2024), we evaluate all algorithms using two key metrics: perplexity and downstream task performance. For Gemma models, following Frantar et al. (2022), we calculate perplexity on C4’s (Raffel et al., 2019) validation set. For PaLM2 models, we calculate perplexity using a $100$ million token subset derived from the PaLM2 training mixture. We use TriviaQA (Joshi et al., 2017), SQuAD (Rajpurkar et al., 2018), NaturalQuestions (Kwiatkowski et al., 2019) and WebQuestions (Berant et al., 2013) to evaluate generation capabilities of the quantized models, and to evaluate their reasoning capabilities, we test on ARC-c, ARC-e (Clark et al., 2018), HellaSwag (Zellers et al., 2019), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020) and WinoGrande (Sakaguchi et al., 2020). For downstream evaluations, we consider the zero-shot setting. Finally, to determine which optimization technique is most effective at solving the layer-wise objective in Equation (1), we evaluate the final solution’s objective value using the same $100$ million tokens that we used to compute perplexity.
@@
Training. All techniques are calibrated using $1280$ data points, where each data point has $2048$ tokens. For OWC, we use a grid size of $50$ to find the most optimal $\gamma$. We used 8 Nvidia H100100100100 GPUs for quantizing the models.
Table 1: Table presents the perplexity evaluations for GPTQ, CD, BCD for INT2 quantization of FFN weights.
Method Epochs Gemma-1 7B Gemma-2 9B Gemma-2 27B PaLM2-Otter PaLM2-Bison
w16a16 $10.384$ $10.683$ $8.682$ $5.984$ $5.298$
w2a16g128
GPTQ - $375.153$ $13.785$ $12.181$ $10.816$ $7.230$
CD $1$ $75.55$ $13.709$ $11.966$ $9.917$ $7.123$
BCD(k=2) $1$ $\mathbf{68.732}$ $\mathbf{13.662}$ $\mathbf{11.873}$ $\mathbf{9.822}$ $\mathbf{7.094}$
@@
Table 2: Table presents the perplexity evaluations for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights. $wx,ay,gz$ in the config column corresponds to $x$-bit weights, $y$-bit activations and group/sub-channel size of $z$.
Config Method Gemma-1 Gemma-2 Gemma-2 PaLM2 PaLM2 PaLM2
7B 9B 27B Gecko Otter Bison
w16a16 - $10.348$ $10.683$ $8.682$ $7.948$ $5.984$ $5.298$
w3a16 OWC $42.885e$ 4 $11.666$ $11.823$ $12.570$ $17.928$ $6.169$
GPTQ $48.157$ $11.382$ $9.681$ $11.347$ $7.176$ $5.774$
CD $19.614$ $11.301$ $9.551$ $10.920$ $7.002$ $5.739$
BCD(k=2) $\mathbf{18.552}$ $\mathbf{11.284}$ $\mathbf{9.526}$ $\mathbf{10.898}$ $\mathbf{6.979}$ $\mathbf{5.733}$
w3a16g128 OWC $21.815$ $11.338$ $9.655$ $11.597$ $8.342$ $5.847$
GPTQ $15.561$ $11.193$ $9.260$ $10.414$ $6.635$ $5.677$
CD $13.496$ $11.182$ $9.237$ $10.273$ $6.655$ $5.656$
BCD(k=2) $13.501$ $11.180$ $9.214$ $10.259$ $6.545$ $5.654$
OWC-CD $14.827$ $11.220$ $9.290$ $10.706$ $6.635$ $5.686$
OWC-CD + CD $13.042$ $11.131$ $\mathbf{9.194}$ $10.143$ $6.528$ $5.650$
OWC-CD + BCD(k=2) $\mathbf{13.004}$ $\mathbf{11.131}$ $9.199$ $\mathbf{10.138}$ $\mathbf{6.527}$ $\mathbf{5.647}$
w4a16 OWC $26.438$ $10.929$ $9.252$ $8.946$ $6.693$ $5.475$
GPTQ $13.355$ $10.896$ $8.923$ $8.764$ $6.249$ $5.417$
CD $12.142$ $\mathbf{10.860}$ $8.910$ $8.694$ $6.195$ $5.407$
BCD(k=2) $\mathbf{12.048}$ $10.863$ $\mathbf{8.899}$ $\mathbf{8.691}$ $\mathbf{6.192}$ $\mathbf{5.405}$
w4a16g128 OWC $11.598$ $10.82$ $8.870$ $8.613$ $6.264$ $5.401$
GPTQ $11.086$ $10.784$ $8.786$ $8.498$ $6.112$ $5.377$
CD $10.838$ $10.777$ $8.781$ $8.456$ $6.097$ $5.373$
BCD(k=2) $10.797$ $\mathbf{10.775}$ $8.786$ $8.454$ $6.097$ $5.372$
OWC-CD $10.981$ $10.786$ $8.802$ $8.519$ $6.106$ $5.377$
OWC-CD + CD $10.766$ $10.776$ $8.778$ $8.436$ $6.092$ $5.371$
OWC-CD + BCD(k=2) $\mathbf{10.760}$ $10.780$ $\mathbf{8.774}$ $\mathbf{8.434}$ $\mathbf{6.091}$ $\mathbf{5.371}$
Results for 2 bit quantization. Table 1 presents the INT2 perplexity numbers for different quantization techniques applied to FFN layers. It can be seen that both our CD and BCD methods have a clear advantage over GPTQ, leading to lower perplexity scores for all models and quantization levels. For example, on PaLM-2 Otter, we see almost $10\%$ improvement in perplexity over GPTQ.
Results for 3,4 bit quantization. Table 2 presents the perplexity numbers for different quantization techniques on 3,4 bit quantization. It can be seen that both our CD and BCD methods have a clear advantage over GPTQ, leading to lower perplexity scores for all models and quantization levels. The difference is more pronounced for lower bit quantization.
Downstream Evals. Downstream evaluations for the PaLM2 and Gemma model families are presented in Tables 6 and 6, respectively. These tables report the average performance across generation and ranking tasks. For a detailed per-task breakdown of the results please refer to Appendix F.2. These results demonstrate that our models achieve comparable, if not superior, performance to GPTQ across all tasks. Notably, our techniques exhibit significant improvements over GPTQ for smaller models like PaLM2-Gecko and Gemma-1 7B.
Layer-wise objective. Finally, in our experiments we also observe that coordinate descent techniques are better at optimizing the layer-wise objective in Equation (1), than GPTQ. For instance, for INT3 per-channel quantization of Gemma-2 9B, the average objective value (relative to all 00’s solution) for GPTQ, for the $1^{\text{st}}$ feed-forward layer is $0.1449$, whereas for CD it is $0.1362$, and for BCD(k=$2$) it is $0.1357$ (which translates to $\sim 6\%$ reduction in the objective value).
Table 3: Table presents the perplexity evaluations for INT2 per-channel quantization of FFN weight with QuIP and FrameQuant, where, GPTQ, CD and BCD are used as subroutines.
Method Gemma-1 7B Gemma-2 9B Gemma-2 27B
w16a16 $10.384$ $10.683$ $8.682$
QuIP GPTQ $25.941$ $13.695$ $11.909$
CD $24.025$ $13.172$ $\mathbf{11.479}$
BCD(k=2) $\mathbf{19.167}$ $\mathbf{13.144}$ $11.547$
FrameQuant GPTQ $26.262$ $12.941$ $10.958$
CD $19.726$ $12.748$ $10.785$
BCD(k=2) $\mathbf{18.242}$ $\mathbf{12.674}$ $\mathbf{10.608}$
Table 4: Table presents the perplexity evaluations for GPTQ, CD, BCD for INT3 quantization of FFN layers with AWQ.
Method Gemma-1 7B Gemma-2 9B PaLM2-Otter
w3a16 w3a16g128 w3a16 w3a16g128 w3a16 w3a16g128
GPTQ + AWQ $25.695$ $14.539$ $11.331$ $11.158$ $7.381$ $6.519$
CD + AWQ $19.703$ $13.254$ $\mathbf{11.255}$ $11.123$ $7.244$ $6.475$
BCD(k=2) + AWQ $\mathbf{18.137}$ $\mathbf{13.205}$ $11.265$ $\mathbf{11.119}$ $\mathbf{7.211}$ $\mathbf{6.474}$
@@
Table 5: Table presents downstream evaluation (zero-shot) numbers for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights for the PaLM2 family of models. Gen, Rank columns correspond to generation and ranking tasks.
@@
w16a16 - $20.1$ $63.02$ $43.84$ $36.23$ $79.51$ $58.57$ $44.29$ $85.44$ $64.55$
w3a16 OWC $15.45$ $55.73$ $39.62$ $15.06$ $59.29$ $41.60$ $39.58$ $75.45$ $61.10$
GPTQ $13.08$ $56.92$ $39.39$ $31.80$ $71.94$ $55.88$ $41.42$ $76.70$ $62.58$
CD $15.82$ $58.19$ $41.24$ $32.34$ $72.03$ $56.16$ $41.66$ $76.71$ $62.69$
BCD(k=2) $16.16$ $57.85$ $41.17$ $32.04$ $72.22$ $56.15$ $41.37$ $76.78$ $62.62$
w3a16g128 OWC $18.54$ $56.95$ $41.59$ $26.62$ $68.09$ $51.50$ $41.54$ $76.51$ $62.52$
GPTQ $17.96$ $58.30$ $42.16$ $33.43$ $72.98$ $57.16$ $43.47$ $76.65$ $63.38$
CD $16.81$ $58.10$ $41.58$ $32.73$ $72.62$ $56.66$ $42.95$ $77.03$ $63.4$
BCD(k=2) $16.33$ $57.81$ $41.22$ $32.13$ $72.67$ $56.46$ $42.67$ $76.85$ $63.18$
OWC-CD $16.73$ $57.74$ $41.34$ $32.90$ $72.40$ $56.60$ $42.23$ $77.02$ $63.11$
OWC-CD + CD $16.37$ $58.22$ $41.48$ $33.53$ $72.64$ $56.99$ $43.09$ $76.91$ $63.38$
w4a16 OWC $16.07$ $59.42$ $42.08$ $33.76$ $72.22$ $56.83$ $43.59$ $77.13$ $63.71$
GPTQ $18.77$ $59.5$ $43.21$ $34.03$ $73.53$ $57.73$ $44.16$ $77.54$ $64.19$
CD $18.20$ $59.95$ $43.25$ $35.01$ $73.42$ $58.06$ $43.97$ $77.55$ $64.12$
BCD(k=2) $17.89$ $59.64$ $42.94$ $35.10$ $73.31$ $58.02$ $43.95$ $77.77$ $64.24$
w4a16g128 OWC $19.00$ $60.42$ $43.85$ $34.52$ $72.83$ $57.5$ $44.45$ $77.71$ $64.41$
GPTQ $19.07$ $60.40$ $43.87$ $35.34$ $73.60$ $58.29$ $44.25$ $77.65$ $64.29$
CD $20.35$ $60.94$ $44.70$ $36.20$ $73.67$ $58.68$ $44.21$ $77.72$ $64.31$
BCD(k=2) $20.44$ $60.69$ $44.59$ $36.11$ $73.54$ $58.57$ $44.17$ $77.56$ $64.2$
OWC-CD $19.80$ $60.39$ $44.15$ $36.33$ $73.62$ $58.71$ $44.75$ $77.70$ $64.52$
OWC-CD + CD $20.47$ $60.85$ $44.70$ $35.98$ $73.59$ $58.54$ $44.77$ $77.52$ $64.42$
Table 6: Table presents downstream evaluation (zero-shot) numbers for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights for both, Gemma-1 and Gemma-2 models. Gen, Rank columns correspond to generation and ranking tasks.
Config Method Gemma-1 7B Gemma-2 9B Gemma-2 27B
Gen. Rank Avg. Gen. Rank Avg. Gen. Rank Avg.
w16a16 $40.6$ $76.96$ $58.35$ $49.34$ $82.6$ $64.36$ $52.57$ $77.53$ $67.55$
w316 OWC $0.76$ $43.95$ $26.6$ $45.84$ $73.51$ $62.44$ $47.29$ $75.21$ $64.04$
GPTQ $19.58$ $58.33$ $40.87$ $45.82$ $73.11$ $62.19$ $49.91$ $75.53$ $65.28$
CD $27.83$ $69.75$ $50.2$ $46.48$ $73.4$ $62.63$ $50.09$ $75.98$ $65.62$
BCD(k=2) $28.21$ $70.08$ $50.51$ $46.23$ $73.34$ $62.49$ $50.34$ $76.03$ $65.75$
w3a16g128 OWC $28.47$ $66.85$ $48.65$ $46.9$ $73.35$ $62.77$ $50.77$ $76.81$ $66.39$
GPTQ $32.94$ $73.23$ $53.82$ $46.62$ $73.52$ $62.76$ $51.26$ $76.49$ $66.4$
CD $34.22$ $73.03$ $54.08$ $46.59$ $73.34$ $62.64$ $51.32$ $77.1$ $66.78$
BCD(k=2) $34.52$ $73.82$ $54.65$ $46.65$ $73.35$ $62.67$ $51.22$ $77.22$ $66.82$
OWC-CD $32.76$ $72.61$ $53.39$ $46.64$ $73.03$ $62.48$ $50.99$ $76.97$ $66.57$
OWC-CD + CD $34.3$ $73.8$ $54.57$ $46.7$ $73.46$ $62.75$ $51.55$ $77.07$ $66.86$
OWC-CD + BCD(k=2) $34.59$ $74.21$ $54.9$ $46.7$ $73.47$ $62.76$ $51.49$ $77.12$ $66.87$
w4a16 OWC $30.69$ $68.82$ $50.5$ $48.49$ $74.1$ $63.85$ $51.63$ $76.9$ $66.79$
GPTQ $35.95$ $74.94$ $55.75$ $48.5$ $74.39$ $64.03$ $52.12$ $76.98$ $67.04$
CD $37.95$ $75.6$ $56.74$ $48.74$ $74.27$ $64.05$ $51.85$ $77.2$ $67.06$
BCD(k=2) $38.11$ $75.95$ $57$ $48.72$ $74.3$ $64.06$ $51.9$ $77.11$ $67.03$
w4a16g128 OWC $38.51$ $75.35$ $56.76$ $49.22$ $73.98$ $64.08$ $52.05$ $77.6$ $67.38$
GPTQ $39.85$ $76.68$ $57.96$ $48.9$ $73.98$ $63.95$ $52.12$ $77.63$ $67.43$
CD $40.01$ $76.79$ $58.08$ $48.9$ $74.05$ $63.99$ $52.29$ $77.6$ $67.47$
BCD(k=2) $40.14$ $76.61$ $58.01$ $49.02$ $73.98$ $64$ $52.22$ $77.65$ $67.48$
OWC-CD $39.44$ $75.99$ $57.43$ $49.26$ $74.21$ $64.23$ $52.15$ $77.75$ $67.51$
OWC-CD + CD $40.19$ $76.89$ $58.19$ $49.06$ $74.13$ $64.1$ $52.24$ $77.71$ $67.52$
OWC-CD + BCD(k=2) $39.94$ $76.52$ $57.89$ $49.1$ $74.17$ $64.14$ $52.14$ $77.69$ $67.47$
4.2 Versatility of CDQuant